OpenStudy (anonymous):

x^2+y^2+3x-4y+4=0 How can I get the radius?

OpenStudy (anonymous):

rewrite as \[(x-h)^2+(y-k)^2=r^2\] by completing the square, and then you will have the radius

OpenStudy (anonymous):

I don't know how to?

OpenStudy (raden):

alternative : if given the equation of circle : x^2+y^2+Ax+By+C=0 so, to get the value of r (radius), use the formula :